# Hubble Tension or Distance Ladder Crisis?

### v2

@@ Abstract
We present an up-to-date compilation of published Hubble constant ($H_{0}$) measurements that are independent of the CMB sound horizon scale. This compilation is categorized into two distinct groups: A. Distance Ladder Measurements: This group comprises 19 recent measurements, primarily from the past four years, utilizing various rung 2 calibrators and rung 3 cosmic distance indicators. B.One-Step Measurements: This category includes 31 measurements of $H_{0}$ that are independent of both the CMB sound horizon scale and the distance ladder approach. These are derived from diverse probes such as Cosmic Chronometers, gamma-ray attenuation, time-delay strong lensing, megamasers etc. Statistical analysis reveals a significant distinction between these two groups. The distance ladder-based measurements yield a weighted mean $H_{0}=72.76\pm 0.50$ km s-1 Mpc-1, while the one-step measurements result in a weighted mean $H_{0}=69.01\pm 0.49$ km s-1 Mpc-1. To quantify the statistical significance of this difference, we implemented a Kolmogorov-Smirnov (KS) test, which yielded a p-value of 0.000102. This extremely low p-value strongly indicates that the two samples are fundamentally distinct, with a probability of less than 0.01% that they are drawn from the same underlying distribution. These findings suggest that late-time one-step measurements of $H_{0}$ are more consistent with early-time measurements, indicating that the core of the Hubble tension lies not between early-time and late-time measurements per se, but specifically between distance ladder measurements and all other $H_{0}$ determinations. This discrepancy points to either a systematic effect influencing all distance ladder measurements or a fundamental physics anomaly affecting at least one rung of the distance ladder.
@@ I Introduction
The main goal of the present analysis is to address these questions using an extensive up-to-date compilation of recent measurements of $H_{0}$ (most published during the past 5 years with emphasis in the last 3 years) which are independent of the sound horizon scale and include both distance ladder measurements and one step measurements that involve no rungs. We thus split the measurements of $H_{0}$ that are independent of the sound horizon scale in two groups: 1. Distance ladder rung based measurements 2. One step measurements that involve direct distance or $H(z)$ probes. We then test the statistical properties of the $H_{0}$ measurements in each group in order to address the above questions. This is the first time that such a statistical meta-analysis is attempted since previous corresponding studies have not separated the two groups of measurements and have also included many early outdated measurementsWang et al. (2023); Lopez-Corredoira (2022); Cao and Ratra (2023).
@@ II.1.3 Connecting BAO and the CMB
These methods, grounded in the physics of the early universe, provide robust constraints on cosmological parameters but are sensitive to assumptions about the cosmological model, primarily the composition and evolution of the universe’s energy density. As such, they are crucial in exploring and potentially resolving tensions in the measured values of $H_{0}$ from local and early universe observations.
@@ II.3.2 Megamasers as Standard Rulers
$D=\frac{v^{2}}{\theta^{2}\omega^{2}}$ (31)
@@ II.3.3 Strong Gravitational Lensing Time Delays
  • Sneppen et al. (2023) Sneppen et al. (2023) derived $H_{0}=67.0\pm 3.6$ km s-1 Mpc-1 using gravitational waves and kilonovae.
@@ II.3.6 Standard Sirens and Kilonovae
Standard sirens, a term coined for gravitational waves (GWs) emitted from compact binary mergers, provide a unique method for measuring the Hubble constant, $H_{0}$, directly from the gravitational wave signal without requiring a cosmic distance ladder Abbott et al. (2017); Hotokezaka et al. (2019); Chen et al. (2018); Ando et al. (2013); Bulla et al. (2022); Sneppen et al. (2023). This section details the principles behind using GWs and their electromagnetic counterparts, kilonovae, as cosmological probes, highlighting their potential to resolve the Hubble tension.
@@ II.3.7 Gamma-Ray Attenuation
Using improved EBL models from comprehensive galaxy surveys such as CANDELS Grogin et al. (2011) and gamma-ray attenuation data from telescopes like Fermi-LAT, recent studies Domínguez et al. (2019, 2023); Desai et al. (2019) have derived new optical depths and used these to calculate the CGRH. This approach provides an independent estimate for $H_{0}$, addressing the tension between local and early universe measurements.
@@ III.1.1 Distance Ladder-Based Measurements
  • Lee et al. used F115W (J-band) for JAGB magnitude measurements, consistent with the theoretical basis of the JAGB method Madore:2020.
@@ III.1.2 One Step Distance Ladder-Independent Measurements
1 Reese (2003) Reese (2003) $61\pm 21$ SZ Effect
@@
17 Sneppen (2023) Sneppen et al. (2023) $67.0\pm 3.6$ Gravitational Waves + Kilonovae
18 Liu (2023) Liu et al. (2023) $59.1^{+3.6}_{-3.5}$ Gravitational Waves + Kilonovae
@@
9. Bulla (2022) Bulla et al. (2022): Utilizing multi-messenger observations from gravitational waves, gamma-ray bursts, and kilonovae associated with neutron star mergers, Bulla et al. reported $H_{0}=69.6\pm 5.5$ km/s/Mpc. This technique harnesses data from different astrophysical events linked by the same originating source, providing a unique way to measure cosmic distances and $H_{0}$ without relying on traditional cosmic ladders.
@@
Figure 2: Same as Fig. 1 but the measurements are binned by year. Notice that for most years the one step $H_{0}$ are consistent with the Planck 18 measurement of $H_{0}$. The only one step measurement that is in significant tension with Planck 18Aghanim et al. (2020) is the TDCOSMO. I. strong lensing time delay measurementMillon et al. (2020) which has been shown to suffer from a systematic issue (related to the mass-sheet degeneracy (MSD)) by a later TDCOSMO analysisShajib et al. (2023). However, more recent TDCOSMO measurements are consistent with Planck 18 and other sound horizon based measurements (eg Birrer et al. (2020); Wong et al. (2024)).
@@ III.2 Results: One step methods vs distance ladder methods
    • One-step (Distance Ladder Independent): $H_{0}=69.01\pm 0.49$ km/s/Mpc
@@
The analysis presented here underscores the critical importance of diverse measurement techniques in cosmology. The agreement between one-step measurements and early universe estimates provided by CMB observations (Planck 18) suggests that the resolution to the Hubble tension may lie in refining our understanding of local distance measurements rather than in revising our models of the early universe.
@@ IV Conclusion-Discussion-Future Prospects
    A fundamental change in physics that impacts the relationships among the distance ladder rungsRuchika et al. (2024); Perivolaropoulos and Skara (2022b); Marra and Perivolaropoulos (2021); Alestas et al. (2021b); Perivolaropoulos and Skara (2021); Huang et al. (2024b); Liu et al. (2024b).
@@
    • Examining the behavior of standard candles and standard sirens across different distance scalesGómez-Valent et al. (2024); Paraskevas and Perivolaropoulos (2023); Dainotti et al. (2023b).
@@
    Theoretical Model Construction: Developing theoretical models that predict localized changes in physics, especially gravitational physics, could provide a framework for understanding potential discrepanciesPerivolaropoulos and Skara (2022c). These models might include:
    • Modified gravity theories with scale-dependent effects.
@@
    • Theories proposing variations in fundamental constants over cosmic time or distancePerivolaropoulos and Skara (2022c).
@@
    • Examining the impact of cosmic variance on local $H_{0}$ measurements.

### v3

@@ Abstract
We present an up-to-date compilation of published Hubble constant ($H_{0}$) measurements that are independent of the CMB sound horizon scale. This compilation is split in two distinct groups: A. Distance Ladder Measurements sample comprising of 19 recent measurements, primarily from the past four years, utilizing various rung 2 calibrators and rung 3 cosmic distance indicators. B.One-Step Measurements sample including 31 measurements of $H_{0}$ that are independent of both the CMB sound horizon scale and the distance ladder approach. These 31 measurements are derived from diverse probes such as Cosmic Chronometers, gamma-ray attenuation, strong lensing, megamasers etc. Statistical analysis reveals a significant distinction between the two samples. The distance ladder-based sample yields a best fit $H_{0}=72.76\pm 0.50$ km s-1 Mpc-1 with $\chi^{2}/dof=0.52$ indicating some correlations. The one-step measurements result in $H_{0}=69.01\pm 0.49$ km s-1 Mpc-1 with $\chi^{2}/dof=1.43$ indicating some internal tension. If two outlier measurements are removed (TDCOSMO.I-2019 known to have systematics and MCP-2020) the best fit of the one step sample reduces to $H_{0}=68.3\pm 0.5$ km s-1 Mpc-1 with $\chi^{2}/dof=1.01$, fully self-consistent and consistent with sound horizon based measurements. A Kolmogorov-Smirnov test yields a p-value of 0.0001 suggesting that the two samples are fundamentally distinct, with a probability of less than 0.01% that they are drawn from the same underlying distribution. These findings suggest that the core of the Hubble tension lies not between early and late-time measurements, but between distance ladder measurements and all other $H_{0}$ determinations. This discrepancy points to either a systematic effect influencing all distance ladder measurements or a fundamental physics anomaly affecting at least one rung of the distance ladder.
@@ I Introduction
The main goal of the present analysis is to address these questions using an extensive up-to-date compilation of recent measurements of $H_{0}$ (most published during the past 5 years with emphasis in the last 3 years) which are independent of the sound horizon scale and include both distance ladder measurements and one step measurements that involve no rungs. We thus split the measurements of $H_{0}$ that are independent of the sound horizon scale in two groups: 1. Distance ladder rung based measurements 2. One step measurements that involve direct distance or $H(z)$ probes. We then test the statistical properties of the $H_{0}$ measurements in each group in order to address the above questions. This is the first time that such a statistical meta-analysis is attempted since previous corresponding studies have not separated the two groups of measurements and have also included many early outdated measurementsWang et al. (2023); Lopez-Corredoira (2022); Cao and Ratra (2023); Verde et al. (2023).
@@ II.1.3 Connecting BAO and the CMB
These methods, grounded in the physics of the early universe, provide robust constraints on cosmological parameters but are sensitive to assumptions about the cosmological model, primarily the composition and evolution of the universe’s energy density. As such, they are crucial in exploring and potentially resolving tensions in the measured values of $H_{0}$ observations.
@@ II.3.2 Megamasers as Standard Rulers
$D=\frac{v}{\theta\omega}$ (31)
@@ II.3.6 Standard Sirens and Kilonovae
Standard sirens, a term coined for gravitational waves (GWs) emitted from compact binary mergers, provide a unique method for measuring the Hubble constant, $H_{0}$, directly from the gravitational wave signal without requiring a cosmic distance ladder Abbott et al. (2017); Hotokezaka et al. (2019); Chen et al. (2018); Ando et al. (2013); Bulla et al. (2022). This section details the principles behind using GWs and their electromagnetic counterparts, kilonovae, as cosmological probes, highlighting their potential to resolve the Hubble tension.
@@ II.3.7 Gamma-Ray Attenuation
Using improved EBL models from comprehensive galaxy surveys such as CANDELS Grogin et al. (2011) and gamma-ray attenuation data from telescopes like Fermi-LAT, recent studies Domínguez et al. (2019, 2023); Desai et al. (2019) have derived new optical depths and used these to calculate the CGRH. This approach provides an one-step sound horizon free estimate for $H_{0}$.
@@ III.1.1 Distance Ladder-Based Measurements
  • Lee et al. used F115W (J-band) for JAGB magnitude measurements, consistent with the theoretical basis of the JAGB method Madore and Freedman (2020).
@@ III.1.2 One Step Distance Ladder-Independent Measurements
1 Reese (2003) Reese (2003) $61\pm 18$ SZ Effect
@@
17 Sneppen (2023) Sneppen et al. (2023) $67.0\pm 3.6$ EPM + Kilonovae
18 Liu (2023) Liu et al. (2023) $59.1^{+3.6}_{-3.5}$ Cluster Lensed Quasar
@@
9. Bulla (2022) Bulla et al. (2022): Utilizing multi-messenger observations from gravitational waves, gamma-ray bursts, and kilonovae associated with neutron star mergers, Bulla et al. reported $H_{0}=69.6\pm 5.5$ km/s/Mpc (weighted mean of the measurements shown in Table 1 of Bulla et al. (2022) using a typical uncertainty from the Table). This technique harnesses data from different astrophysical events linked by the same originating source, providing a unique way to measure cosmic distances and $H_{0}$ without relying on traditional cosmic ladders.
@@
Figure 2: Same as Fig. 1 but the measurements are binned by year. Notice that for most years the one step $H_{0}$ are consistent with the Planck 18 measurement of $H_{0}$. The only one step measurement that is in significant tension with Planck 18Aghanim et al. (2020) is the TDCOSMO. I. strong lensing time delay measurementMillon et al. (2020) which has been shown to suffer from a systematic issue (related to the mass-sheet degeneracy (MSD)) by later TDCOSMO analysesShajib et al. (2023); Birrer et al. (2020). However, more recent TDCOSMO measurements are consistent with Planck 18 and other sound horizon based measurements (eg Birrer et al. (2020); Wong et al. (2024)).
@@ III.2 Results: One step methods vs distance ladder methods
To combine the multiple measurements of $H_{0}$ from both distance ladder dependent and independent methods, we employed a weighted average approach. This method assumes that all measurements are independent of each other. The procedure is as follows:
  1. 1.
    Weighted Mean Calculation: Each $H_{0}$ measurement is weighted by the inverse square of its uncertainty, giving more weight to measurements with smaller uncertainties. The weighted mean $H_{0}$ is calculated as:
    $\overline{H_{0}}=\frac{\sum_{i}H_{0,i}/\sigma_{i}^{2}}{\sum_{i}1/\sigma_{i}^{2}}$ (60)
    where $H_{0,i}$ is the $i$-th measurement and $\sigma_{i}$ is its uncertainty.
  2. 2.
    Uncertainty Calculation: The uncertainty of the weighted mean is estimated as the inverse square root of the sum of weights:
    $\sigma_{\overline{H_{0}}}=\frac{1}{\sqrt{\sum_{i}1/\sigma_{i}^{2}}}$ (61)
This method assumes that all measurements are independent of each other, allowing us to treat each measurement as providing separate information. However, it’s important to note that in reality, some measurements might have correlations due to shared data or methodologies. The independence assumption, while commonly used for simplicity, may lead to an underestimation of the final uncertainty if significant correlations exist between measurements.
@@
    • One-step (Distance Ladder Independent): $H_{0}=69.01\pm 0.49$ km/s/Mpc. This value decreases further to $H_{0}=68.3\pm 0.5$ km/s/Mpc if two outliers are removed from the sample as discussed below.
@@
III.2.1 $\chi^{2}$ analysis and robustness of results
To further assess the consistency and reliability of our findings, we conducted a comprehensive $\chi^{2}$ analysis on various subsets of the distance ladder independent dataset. This analysis serves to evaluate the internal consistency of different measurement methods and identify potential sources of tension within the data.
We constructed and minimized the $\chi^{2}$ statistic for each subset, defined as:
$\chi^{2}=\sum_{i}\frac{(H_{0,i}-H_{0,\text{best fit}})^{2}}{\sigma_{i}^{2}}$ (62)
where $H_{0,i}$ are the individual measurements, $\sigma_{i}$ their uncertainties, and $H_{0,\text{best fit}}$ the best-fit value that minimizes $\chi^{2}$.
Table 3 presents the results of this analysis for various subsets of the data:
Table 3: Analysis of Various Subsets of the One Step - Sound Horizon and Distance Ladder Independent Sample. The data numbers correspond to Table 2. Notice the dramatic improvement of the self consistency of the full sample (decrease of the reduced $\chi^{2}_{\text{red}}\equiv\chi^{2}/dof$ ) if the two outliers (TDCOSM.I and MCP-SH0ES) are removed.
Dataset $H_{0}$ (km s-1 Mpc-1) $\chi^{2}_{\text{min}}$ $\chi^{2}_{\text{red}}$ DoF Data points
Full Sample $69.0\pm 0.5$ 43.14 1.44 30 All
Full Sample (excl. 3,4) $68.3\pm 0.5$ 28.16 1.01 28 All except 3,4
GW and Kilonovae EPM $66.3\pm 2.0$ 8.13 1.16 7 6,9,14,15,17,18,22,27
Lensing All $71.4\pm 1.0$ 12.00 1.33 9 3,5,11,12,19,20,26,29-31
Lensing (no TDCOSMO I) $69.7\pm 1.2$ 7.14 0.89 8 5,11,12,19,20,26,29-31
Cosmic Chronometers $68.3\pm 2.0$ 1.80 0.90 2 8,10,13
Megamasers $72.3\pm 2.7$ 1.39 1.39 1 2,4
Our analysis reveals that most subsamples are self-consistent and in agreement with each other, as evidenced by their reduced $\chi^{2}$ values close to 1. However, we observe some tension in the full dataset (reduced $\chi^{2}=1.43$) and in two subsamples: Megamasers and Lensing All.
The tension in the full dataset and the in the above two subsamples can be attributed to two notable outliers: data points 3 and 4 of Table 2, associated with TDCOSMO I lensingMillon et al. (2020) and the Megamasers MCP+SH0ES Pesce et al. (2020) points respectively. The TDCOSMO I measurement (point 3) has been shown to suffer from systematic issues related to the mass-sheet degeneracy (MSD), as demonstrated by subsequent TDCOSMO analysesShajib et al. (2023); Birrer et al. (2020). When we exclude these two outliers (Full Data excl. 3,4), the reduced $\chi^{2}$ ($\chi^{2}$ per dof) improves significantly to 1.00585, indicating excellent consistency among the remaining data points. In addition, the consistency of the remaining one step-sound horizon free measurements with sound horizon based measurements improves further and the new best fit value of $H_{0}$ becomes $H_{0}=68.3\pm 0.5$km s-1 Mpc-1.
The GW+kilonovae subset combines measurements based on gravitational waves with measurements based on kilonovae. Even though this subset encompasses diverse methods, it shows good internal consistency (reduced $\chi^{2}=1.16106$) and yields a best-fit $H_{0}$ value of $66.3\pm 2.0$ km s-1 Mpc-1, which is in close agreement with the Planck 18 result. This supports the potential of gravitational wave standard sirens as a promising independent probe of $H_{0}$.
The Lensing subset, after excluding the TDCOSMO I outlier, demonstrates improved consistency (reduced $\chi^{2}=0.892011$) and a best-fit $H_{0}$ of $69.7\pm 1.2$ km s-1 Mpc-1, aligning more closely with other independent measurements.
Cosmic Chronometers and the revised Lensing subset both show excellent internal consistency (reduced $\chi^{2}<1$) and yield $H_{0}$ values that are in good agreement with the overall trend of one-step measurements.
The Megamasers subset, while showing some tension (reduced $\chi^{2}=1.38689$), is based on only two data points. The tension is apparently due to the outlier MCP+SH0ES datapointPesce et al. (2020) discussed above.
The outlier data: Here we discuss in some more detail the two outlier one-step $H_{0}$ measurements and point out potential sources of systematic errors in these measurements.
TDCOSMO.I measurement (2019) Millon et al. (2020) $H_{0}=74.2\pm 1.6$
: The TDCOSMO I measurement (point 3 of Table II) suffers from systematic issues related to the mass-sheet degeneracy (MSD). This early analysis by Millon et al. (2020) assumed simple parametric mass models (power-law or composite) for the lens galaxies, which implicitly break the MSD. However, these assumptions can potentially bias the $H_{0}$ measurement or underestimate the errors if the true mass profile deviates from these simple models.
Subsequent analyses, particularly Birrer et al. (2020) (TDCOSMO IV) and Shajib et al. (2023), have effectively replaced this measurement by addressing the MSD issue more rigorously. These later studies have much larger uncertainties because they relax the assumptions on the mass profile and instead use stellar kinematics to constrain the MSD. Ref. Birrer et al. (2020) (measurment 5 of Table 2) used single-aperture stellar kinematics for the TDCOSMO sample, while Shajib et al. (2023) (measurment 19 of Table 2) used spatially-resolved kinematics for RXJ1131--1231. By allowing more freedom in the mass model and constraining it with kinematic data, these analyses provide a more robust, albeit less precise, measurement of $H_{0}$. The increased uncertainties reflect the additional degeneracies introduced when the mass profile is not assumed to follow a simple parametric form.
MCP-SH0ES measurement (2020) Pesce et al. (2020) $H_{0}=73.9\pm 3.0$
: This analysis attempts to estimate $H_{0}$ using 6 megamasers in galaxies with relatively low redshifts ($z\approx 0.002-0.034$), which are barely in the Hubble flow. Given these low redshifts, it is expected that peculiar velocities should lead to increased uncertainties in the $H_{0}$ measurement. However, the reported uncertainty of ±$\pm 3.0$ km s-1 Mpc-1 seems surprisingly small given this effect. Notably, this measurement is in tension with previous megamaser-based $H_{0}$ determinations, which found significantly lower values (e.g., $H_{0}=66.0\pm 6.0$ km s-1 Mpc-1 from Gao et al. 2016 Gao et al. (2016) and $H_{0}=68\pm 9$ km s-1 Mpc-1 from Kuo et al. 2013 Kuo et al. (2013)). The lack of any updated megamaser-based $H_{0}$ measurements since 2020 raises further questions about the robustness of this result. Other potential sources of systematics include the assumptions made in the disk modeling, particularly the treatment of warps and eccentricity, as well as possible biases in the selection of megamaser galaxies (6 chosen out of more than 15 knownGao et al. (2017)). The reliance of the analysis on external galaxy flow models to correct for peculiar velocities introduces additional model-dependent uncertainties. Furthermore, the small sample size makes the measurement susceptible to cosmic variance effects.
These arguments along with the results of Table 3 support the robustness of our findings across various subsets of the distance ladder independent data. The consistency observed after removing known problematic outliers strengthens our conclusion that one-step methods tend to favor lower $H_{0}$ values, in better agreement with Planck 18 results, compared to distance ladder-dependent measurements.
The distance ladder sample: Correlations and over fitting
Interestingly, the analysis of the full distance ladder dependent dataset yields a notably different result. With a best-fit $H_{0}$ of $72.7^{+0.5}_{-0.5}$ km s-1 Mpc-1, this dataset shows a clear preference for higher $H_{0}$ values compared to most of the independent subsets. However, the most striking feature is the unusually low reduced $\chi^{2}$ value of 0.52.
A reduced $\chi^{2}$ significantly below 1, as observed in the distance ladder dependent dataset, suggests that the model is ”overfitting” the data. This could be due to several factors:
  1. 1.
    Overestimated uncertainties: If the reported uncertainties in the distance ladder measurements are consistently larger than the true uncertainties, it would lead to an artificially low $\chi^{2}$ value.
  2. 2.
    Correlation between measurements: If there are unaccounted correlations between different distance ladder measurements, it could lead to an underestimation of the true variance in the data, resulting in a lower $\chi^{2}$.
  3. 3.
    Publication bias: There might be a tendency to publish results that are in agreement with previous measurements, leading to an artificially low scatter in the published values.
This unusually low reduced $\chi^{2}$ in the distance ladder dependent dataset contrasts sharply with the more typical values seen in the independent dataset subsets. It suggests that the uncertainties in the distance ladder measurements may be overestimated, or that there are significant correlations or dependencies among these measurements that are not being accounted for in the analysis.
Interpretation and the Road Ahead:
This analysis highlights the importance of critically examining individual measurements and their potential systematic effects. It also demonstrates the value of using multiple independent methods to constrain $H_{0}$, as this approach allows for the identification and mitigation of method-specific biases, systematics or new physics.
@@ III.2.1 $\chi^{2}$ analysis and robustness of results
The analysis presented here underscores the critical importance of diverse measurement techniques in cosmology. The agreement between one-step measurements and sound horizon based estimates provided by CMB observations (Planck 18) suggests that the resolution to the Hubble tension may lie in refining our understanding of local distance ladder measurements rather than in revising our models of the early universe.
@@ IV Conclusion-Discussion-Future Prospects
    A fundamental change in physics that impacts the relationships among the distance ladder rungsRuchika et al. (2024); Perivolaropoulos and Skara (2022b); Marra and Perivolaropoulos (2021); Alestas et al. (2021b); Perivolaropoulos and Skara (2021); Huang et al. (2024b); Liu et al. (2024b); Desmond and Sakstein (2020); Desmond et al. (2019).
@@
    • Examining the behavior of standard candles and standard sirens across different distance scalesGómez-Valent et al. (2024); Paraskevas and Perivolaropoulos (2023); Dainotti et al. (2023b); Perivolaropoulos and Skara (2022b, 2021).
@@
    Theoretical Model Construction: Developing theoretical models that predict localized changes in physics, especially gravitational physics, could provide a framework for understanding potential discrepanciesPerivolaropoulos and Skara (2022c); Desmond and Sakstein (2020); Desmond et al. (2019). These models might include:
    • Modified gravity theories with scale-dependent effects Perivolaropoulos and Skara (2022c); Desmond and Sakstein (2020); Desmond et al. (2019).
@@
    • Theories proposing variations in fundamental constants over cosmic time or distance.
@@
    • Examining the impact of cosmic variance on distance ladder $H_{0}$ measurements.